# CasualHDRSplat: Robust High Dynamic Range 3D Gaussian Splatting from Casually Captured Videos

### v1

@@ 6 Appendix
6.1 HDR editing
Figure 7: Qualitative results of HDR editing with various exposure times.
The results in Figure 7 demonstrate that our method can accurately learn HDR scenes and the brightness of the rendered images can be adjusted by manually changing the exposure time.
6.2 More experimental results of exposure time estimation.
Table 7: Quantitative results for exposure time estimation.
@@ 6.1 HDR editing
where $f$ is the focal length, $t$ is the exposure time, $g$ is the gain, $a$ is the aperture f-number. Like most 3D HDR methods[17, 4, 13], the $f$, $g$ and $a$ are fixed but unknown in experiments. Thus, the exposure times that we estimate have a constant scale with the ground truth exposure times. To evaluate the exposure time estimation, we used metrics such as Pearson correlation coefficient, Spearman rank correlation coefficient, and Kendall Tau-b correlation coefficient. A correlation value greater than 0.7 is generally considered a strong correlation. The quantitative results are shown in Table 7, the results show that we can estimate exposure times :
Figure 9 shows the comparison between the estimated exposure times and the ground truth exposure times for each training image in scenes Toufu-vicon (above) and Girls-vicon (bottom). The results are scaled uniformly, and it can be observed that the estimated exposure times closely follow the trend of the ground truth exposure times.
6.3 More experimental results under training view.
Figure 8: Qualitative comparison on synthetic dataset (Trolley, Factory, Cozyroom) under training view. BAD-Gaussians can deblur the training views as ours. However, due to the failure of pose optimization in the BAD-Gaussians, its image are misaligned with others.
The results in Figure 8 demonstrate that our method effectively deblurs images under training views and achieves better image quality compared to other methods. It is worth noting that while BAD-Gaussians [64] is also capable of deblurring images under training view, its lack of robustness to varying brightness conditions leads to pose optimization failure. As a result, The performance of deblurring is poor, even causing misalignment in the images under the training view.
Figure 9: Comparison between estimated exposure times and ground truth exposure times. Toufu-vicon (above) and Girls-vicon (bottom).
Table 8: Quantitative comparisons on the synthetic datasets in term of deblur.
@@
BAD-GS [64] 24.32 0.73 0.12 25.87 0.79 0.23 19.06 0.62 0.19 23.37 0.79 0.11
BAD-GS+bilagrid [56] 28.25 0.79 0.08 31.99 0.86 0.06 22.16 0.65 0.15 26.48 0.81 0.09
CasualHDRSplat-random (ours) 31.20 0.88 0.05 32.95 0.87 0.10 23.65 0.69 0.12 29.60 0.84 0.05
CasualHDRSplat-gt (ours) 32.00 0.91 0.07 34.53 0.96 0.05 29.35 0.87 0.08 33.01 0.93 0.04
Table 9: Quantitative comparisons on the ScanNet dataset in term of deblur (using BRISQUE metric).
BAD-GS [64] 42.78 60.48 57.37 42.72 67.30 65.01 55.94
CasualHDRSplat-random (ours) 38.08 47.65 53.37 37.55 62.35 58.23 49.53
We added comparison against the bilateral grid method [56] applied to gsplat [59] and BAD-Gaussians [64] in Table 8 and Table 13. The bilateral grids applied to NeRFs and 3DGS gives robustness to large appearance changes, enabling high quality 3D LDR reconstruction and mid-tone rendering quality. However, bilateral grids are not compatible with representing a 3D HDR scene, thus gives degraded renderings on high-contrast, over-exposured and under-exposured views.
In addition, we also evaluated the quantitative metrics for deblurring on public real-world datasets, e.g. ScanNet datastes. Due to most images of ScanNet dataset are motion-blurred, we can not find sharp reference images for evaluating, thus we utilize the no-reference image quality metric BRISQUE [38] to quantitatively compare the deblurring performance between our method and BAD-Gaussians [64], as shown in Table 9.
Table 10: Quantitative comparisons on Realsense and SmartPhone dataset under novel view.
@@
gsplat [59] 25.04 0.83 0.27 29.88 0.81 0.24
BAD-Gaussians [64] 23.31 0.78 0.28 30.05 0.82 0.24
HDR-Plenoxels [17] 27.13 0.81 0.33 30.91 0.82 0.29
Gaussian-W [62] 27.57 0.84 0.28 30.89 0.83 0.26
CasualHDRSplat-random (ours) 28.56 0.84 0.22 32.75 0.87 0.17
CasualHDRSplat-gt (ours) 29.19 0.87 0.16 32.84 0.91 0.18
@@
gsplat [59] 23.20 0.82 0.16 25.99 0.81 0.11 24.34 0.81 0.28 23.81 0.77 0.28
gsplat+bilagrid [56] 25.26 0.78 0.14 25.47 0.77 0.16 30.48 0.82 0.17 26.76 0.69 0.25
BAD-GS [64] 24.28 0.78 0.14 26.93 0.82 0.11 24.22 0.82 0.24 23.95 0.77 0.28
BAD-GS+bilagrid [56] 25.12 0.77 0.17 25.63 0.77 0.15 30.52 0.83 0.17 26.18 0.71 0.23
HDR-Plenoxels [17] 19.39 0.53 0.65 26.87 0.81 0.15 17.90 0.51 0.69 26.73 0.84 0.30
Gaussian-W [62] 26.13 0.83 0.15 27.99 0.82 0.11 26.38 0.83 0.29 26.88 0.86 0.25
CasualHDRSplat-random (ours) 28.30 0.83 0.13 28.79 0.83 0.09 30.87 0.90 0.15 32.00 0.90 0.19
CasualHDRSplat-gt (ours) 30.81 0.87 0.12 29.71 0.85 0.08 31.34 0.92 0.12 32.39 0.91 0.17
Figure 10: Qualitative comparison for pose estimation on the Girls-vicon sequence of the Realsense dataset.
Figure 11: Qualitative comparison for pose estimation on the Toufu-vicon sequence of the Realsense dataset.
6.4 More experimental results under novel view.
@@
Figure 12: Qualitative comparison with bilateral method on Smartphone dataset under novel view. It is better to view the results on a monitor with high resolution and a gamut coverage close or better than sRGB.
In addition, we compare against the bilateral grid method [56] applied to gsplat [59] and BAD-Gaussians [64], as shown in Table 10 and Figure 12. As aforementioned, with bilateral grid [56], BAD-Gaussians [64] cannot represent the HDR details of the 3D scenes, thus yields degraded renderings in the high-contrast areas. As it is showed in the Figure 12, the girls in the Fish sequence of the Smartphone dataset are over-exposed in some views, thus exhibits under-saturation; Meanwhile, the duck in the Building sequence of the Smartphone dataset has lost its details and exhibits artifacts on its edge.
6.5 Video results of CasualHDRSplat
@@
Figure 13: Qualitative comparison on synthetic dataset(Cozyroom, Factory, Outdoorpool) under novel view.
Figure 14: Qualitative comparison on Smartphone dataset under novel view.
6.6 More experimental results on pose estimation.

### v2

@@ Appendix A Appendix
A.1. More experimental results of exposure time estimation.
Table 8. Quantitative results for exposure time estimation.
@@
where $f$ is the focal length, $t$ is the exposure time, $g$ is the gain, $a$ is the aperture f-number. Like most 3D HDR methods(Jun-Seong et al., 2022; Cai et al., 2024; Huang et al., 2022), the $f$, $g$ and $a$ are fixed but unknown in experiments. Thus, the exposure times that we estimate have a constant scale with the ground truth exposure times. To evaluate the exposure time estimation, we used metrics such as Pearson correlation coefficient, Spearman rank correlation coefficient, and Kendall Tau-b correlation coefficient. A correlation value greater than 0.7 is generally considered a strong correlation. The quantitative results are shown in Table 8, the results show that we can estimate exposure times :
Figure 9 shows the comparison between the jointly-optimized exposure times and the ground truth exposure times for each training image in scenes Toufu-vicon (left) and Girls-vicon (right). The results are scaled uniformly, and it can be observed that the estimated exposure times closely follow the trend of the ground truth exposure times.
A.2. More experimental results under training view.
Figure 8. Qualitative comparison on synthetic dataset (Trolley, Factory, Cozyroom) under training view. BAD-Gaussians can deblur the training views as ours. However, due to the failure of pose optimization in the BAD-Gaussians, its image are misaligned with others.
The results in Figure 8 demonstrate that our method effectively deblurs images under training views and achieves better image quality compared to other methods. It is worth noting that while BAD-Gaussians (Zhao et al., 2024) is also capable of deblurring images under training view, its lack of robustness to varying brightness conditions leads to pose optimization failure. As a result, The performance of deblurring is poor, even causing misalignment in the images under the training view.
Figure 9. Comparison between estimated exposure times and ground truth exposure times. Girls-vicon (left) and Toufu-vicon (right).
Table 9. Quantitative comparisons on the synthetic datasets in term of deblur.
@@
BAD-Gaussians (Zhao et al., 2024) 24.32 0.73 0.12 25.87 0.79 0.23 19.06 0.62 0.19 23.37 0.79 0.11
BAD-Gaussians+bilagrid (Wang et al., 2024a) 28.25 0.79 0.08 31.99 0.86 0.06 22.16 0.65 0.15 26.48 0.81 0.09
Casual3DHDR-random (ours) 31.20 0.88 0.05 32.95 0.87 0.10 23.65 0.69 0.12 29.60 0.84 0.05
Casual3DHDR-gt (ours) 32.00 0.91 0.07 34.53 0.96 0.05 29.35 0.87 0.08 33.01 0.93 0.04
We added comparison against the bilateral grid method (Wang et al., 2024a) applied to gsplat (Ye et al., 2024) and BAD-Gaussians (Zhao et al., 2024) in Table 9 and Table 13. The bilateral grids applied to NeRFs and 3DGS gives robustness to large appearance changes, enabling high quality 3D LDR reconstruction and mid-tone rendering quality. However, bilateral grids are not compatible with representing a 3D HDR scene, thus gives degraded renderings on high-contrast, over-exposured and under-exposured views.
Table 10. Quantitative comparisons on Realsense and SmartPhone dataset under novel view.
@@
gsplat (Ye et al., 2024) 25.04 0.83 0.27 29.88 0.81 0.24
BAD-Gaussians (Zhao et al., 2024) 23.31 0.78 0.28 30.05 0.82 0.24
HDR-Plenoxels (Jun-Seong et al., 2022) 27.13 0.81 0.33 30.91 0.82 0.29
Gaussian-W (Zhang et al., 2024) 27.57 0.84 0.28 30.89 0.83 0.26
Casual3DHDR-random (ours) 28.56 0.84 0.22 32.75 0.87 0.17
Casual3DHDR-gt (ours) 29.19 0.87 0.16 32.84 0.91 0.18
@@
gsplat (Ye et al., 2024) 23.20 0.82 0.16 25.99 0.81 0.11 24.34 0.81 0.28 23.81 0.77 0.28
gsplat+bilagrid (Wang et al., 2024a) 25.26 0.78 0.14 25.47 0.77 0.16 30.48 0.82 0.17 26.76 0.69 0.25
BAD-Gaussians (Zhao et al., 2024) 24.28 0.78 0.14 26.93 0.82 0.11 24.22 0.82 0.24 23.95 0.77 0.28
BAD-Gaussians+bilagrid (Wang et al., 2024a) 25.12 0.77 0.17 25.63 0.77 0.15 30.52 0.83 0.17 26.18 0.71 0.23
HDR-Plenoxels (Jun-Seong et al., 2022) 19.39 0.53 0.65 26.87 0.81 0.15 17.90 0.51 0.69 26.73 0.84 0.30
Gaussian-W (Zhang et al., 2024) 26.13 0.83 0.15 27.99 0.82 0.11 26.38 0.83 0.29 26.88 0.86 0.25
Casual3DHDR-random (ours) 28.30 0.83 0.13 28.79 0.83 0.09 30.87 0.90 0.15 32.00 0.90 0.19
Casual3DHDR-gt (ours) 30.81 0.87 0.12 29.71 0.85 0.08 31.34 0.92 0.12 32.39 0.91 0.17
Figure 10. Qualitative comparison for pose estimation on the Girls-vicon sequence of the Realsense dataset.
Figure 11. Qualitative comparison for pose estimation on the Toufu-vicon sequence of the Realsense dataset.
A.3. More experimental results under novel view.
@@
Figure 12. Qualitative comparison with bilateral method on Smartphone dataset under novel view. It is better to view the results on a monitor with high resolution and a gamut coverage close or better than sRGB.
In addition, we compare against the bilateral grid method (Wang et al., 2024a) applied to gsplat (Ye et al., 2024) and BAD-Gaussians (Zhao et al., 2024), as shown in Table 10 and Figure 12. As aforementioned, with bilateral grid (Wang et al., 2024a), BAD-Gaussians (Zhao et al., 2024) cannot represent the HDR details of the 3D scenes, thus yields degraded renderings in the high-contrast areas. As it is showed in the Figure 12, the girls in the Fish sequence of the Smartphone dataset are over-exposed in some views, thus exhibits under-saturation; Meanwhile, the duck in the Building sequence of the Smartphone dataset has lost its details and exhibits artifacts on its edge.
Figure 13. Qualitative comparison on synthetic dataset(Cozyroom, Factory, Outdoorpool) under novel view.
Figure 14. Qualitative comparison on Smartphone dataset under novel view.
A.4. More experimental results on pose estimation.
A.5. Video results of Casual3DHDR